OpenStudy (anonymous):

The population of a local species of beetle can v found using an infinite geometric series where a sub(1)=880 and the common ration is 1/4. Write the sum in sigma notation, and calculate the sum (if possible) that will be the upper limit of this population.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

the terms decrease geometrically to a smaller and smaller magnitude, so you CAN approximate the sum.

OpenStudy (solomonzelman):

For \(\large\color{black}{ a_1=880}\) and \(\large\color{black}{ r=¼ }\)

OpenStudy (anonymous):

Is the sum going to be divergent?

OpenStudy (anonymous):

Well, I guess not, because we can approximate an answer...

OpenStudy (solomonzelman):

You will have ∞ on top of the Σ.\(\huge\color{white}{ \rm │ }\) AND the equation (next to Σ) will be something like \(\large\color{black}{880(¼)^n }\) And index (below Σ) will be n=0 .\(\LARGE\color{white}{ \rm │ }\)

OpenStudy (solomonzelman):

Can you write the sigma notation ?

OpenStudy (anonymous):

there are thwo choices where you have n: either i, or i-1

OpenStudy (anonymous):

What do you mean? write out the answer choices?

OpenStudy (solomonzelman):

sigma with imaginary numbers? I can't really see why is that ...

OpenStudy (anonymous):

on the bottom (the index) it says i=1, not n=0

OpenStudy (anonymous):

That's on all of them

OpenStudy (solomonzelman):

Why can't it be \[\Large \sum_{n=0}^{∞} ~880\left(\begin{matrix} \frac{1}{4} \end{matrix}\right)^n\]

OpenStudy (anonymous):

I have no idea- that just not one of the answers...

OpenStudy (solomonzelman):

Ohh i is for index, it is not an imaginary.

OpenStudy (solomonzelman):

\[\Large \sum_{i=0}^{∞} ~880\left(\begin{matrix} \frac{1}{4} \end{matrix}\right)^i\]

OpenStudy (anonymous):

That makes so much more sense!!

OpenStudy (solomonzelman):

but putting an "i" instead of "n" is very confusing.

OpenStudy (solomonzelman):

I never liked "i" as just a variable (that is not an imaginary)

OpenStudy (anonymous):

It really is. maybe not in really early math, when you have no idea what an imaginary number is, but now that the imaginary is part of our mental math vocabulary, it shouldn't be used like that anymore.

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

I hate when people solve equations with i and a bunch of other variables and say solve for x.... (I keep asking "is i an imaginary or...)

OpenStudy (solomonzelman):

Do you have any more questions about your question ?

OpenStudy (anonymous):

I know! It makes math even more confusing that it already is.

OpenStudy (anonymous):

Should the exponent of the parentheses be i or i-1?

OpenStudy (anonymous):

If the index is i=1

OpenStudy (solomonzelman):

it is i-1 if you start from n=1, and it should be just i, if you start from n=0

OpenStudy (anonymous):

Thank you! I'm just starting from i=1 because that's what the problem did. It doesn't make any sense to me.

OpenStudy (solomonzelman):

It is either \[\Large \sum_{i=0}^{∞} 880\left(\begin{matrix} \frac{1}{4}\end{matrix}\right)^i\] or \[\color{reblue}{ \Large \sum_{i=1}^{∞} 880\left(\begin{matrix} \frac{1}{4}\end{matrix}\right)^{i-1}}\]